OpenStudy (anonymous):

x-2y+3z=18 9x+2y-z=-2 -6x-y+2z=4 I need to solve the system for x,y,z

OpenStudy (wolf1728):

A) x-2y+3z=18 B) 9x+2y-z=-2 C) -6x-y+2z=4 A) -2y = -x -3z +18 B) 2y = -9x +z -2 Adding these 2 equations yields equation D D) -10x -2z +16 = 0 B) 9x+2y-z=-2 C) -6x-y+2z=4 Multiplying this equation 2 yields C2) -12x -2y + 4z = 8 Adding these equations together B) 9x+2y-z=-2 C2) -12x -2y + 4z = 8 produces E) -3x + 3z = 6 then we multiply equation D) by 1.5 producing D2) -15x -3z = -24 Adding these 2 together produces -18x = -18 x = 1 That should be a good start for you.

OpenStudy (anonymous):

x - 2y + 3z = 18 9x + 2y - z = -2 --------------add 10x + 2z = 16 9x + 2y - z = -2 -6x - y + 2z = 4 -->(2) --------------- 9x + 2y - z = -2 -12x - 2y + 4z = 8 (result of multiplying by 2) ----------------add -3x + 3z = 6 10x + 2z = 16 -->(-3) -3x + 3z = 6 --> (2) ---------------- -30x - 6z = - 48 (result of multiplying by -3) -6x + 6z = 12 (result of multiplying by 2) ----------------add -36x = - 36 x = 1 now sub 1 in for x -3x + 3z = 6 -3(1) + 3z = 6 -3 + 3z = 6 3z = 6 + 3 3x = 9 z = 3 now sub 1 in for x and 3 in for y in any of the original equations x - 2y + 3z = 18 1 - 2y + 3(3) = 18 1 - 2y + 9 = 18 -2y + 10 = 18 -2y = 18 - 10 -2y = 8 y = -4 check by subbing in known values 9x + 2y - z = -2 9(1) + 2(-4) - 3 = -2 9 - 8 - 3 = -2 9 - 11 = -2 -2 = -2 (correct) x = 1, y = -4, z = 3 or (1,-4,3)